OpenStudy (anonymous):

Prove that for all integers n > 0, 8n – 3n is divisible by 5. check my answer?

OpenStudy (anonymous):

Base Case: n=1 8^1-3^1=5 5/5=1 Induction Hypothesis: 8^k-3^k is divisible by 3 8^(k+1)-3^(k+1) 8^(2)-3^(2) 64-9=53 53/3

jimthompson5910 (jim_thompson5910):

I think you meant to say `Induction Hypothesis: 8^k-3^k is divisible by 5`

jimthompson5910 (jim_thompson5910):

if we assume that claim is true, then we can say 8^k-3^k = 5q where q is some integer. Agreed?

OpenStudy (anonymous):

what is the significance of 5q??

jimthompson5910 (jim_thompson5910):

it just means "5 times some integer"

jimthompson5910 (jim_thompson5910):

if I said "100 is divisible by 5" I can state that as 100 = 5q In this case, q = 20

OpenStudy (anonymous):

ok, yeah. so that just means divisible by 5

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

the goal is to prove that if 8^k-3^k = 5q is assumed to be true, then 8^(k+1)-3^(k+1) = 5s where q and s are integers

jimthompson5910 (jim_thompson5910):

here are 2 hints: 1) isolate the `8^k` in `8^k-3^k = 5q` to get `8^k = 5q + 3^k` 2) break down `8^(k+1)-3^(k+1) ` into `8*8^k-3*3^k`

OpenStudy (anonymous):

i do not understand.8^(k+1) is just 64, no? not 64^K?

jimthompson5910 (jim_thompson5910):

how are you getting 64? you can't replace k with 1

OpenStudy (anonymous):

i didn't know i couldn't do that.

jimthompson5910 (jim_thompson5910):

you can replace k with 1 if you wanted to prove that specific piece. But we're trying to prove the whole thing in general (for any number k)

jimthompson5910 (jim_thompson5910):

do you see how I turned `8^(k+1)-3^(k+1)` into `8*8^k-3*3^k` ?

OpenStudy (anonymous):

no, that was my next question. i thought k+1 would turn into ^2

jimthompson5910 (jim_thompson5910):

recall that \[\LARGE x^{y+z} = x^y * x^z\]

jimthompson5910 (jim_thompson5910):

so \[\Large 8^{k+1} = 8^k*8^1 = 8^k*8 = 8*8^k\] make sense?

OpenStudy (anonymous):

yes. thank you. and now i have to do the same thing for the other side, right?

jimthompson5910 (jim_thompson5910):

for the 3^(k+1) part you mean? yes

OpenStudy (anonymous):

ok, so i just get 3^k*3+5q=6k+5q

jimthompson5910 (jim_thompson5910):

we have 8*8^k-3*3^k now replace the 8^k with 5q + 3^k 8*8^k-3*3^k 8*(5q + 3^k) - 3*3^k 8*5q + 8*3^k - 3*3^k do you see what's next?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

40q + 24^k - 9^k 40q + 15^k

jimthompson5910 (jim_thompson5910):

8*3^k isn't equal to 24^k

jimthompson5910 (jim_thompson5910):

8*5q + 8*3^k - 3*3^k 5*8q + 8*3^k - 3*3^k 5*8q + (8*3^k - 3*3^k) 5*8q + (8 - 3)*3^k 5*8q + 5*3^k 5*(8q + 3^k) 5*s where s = 8q + 3^k and s is an integer. So this all proves that if 8^k - 3^k = 5q, then 8^(k+1)-3^(k+1) = 5s q and s are integers

OpenStudy (anonymous):

im going to have to study that, tomorrow. can you look at a different kind of problem with me?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

1^3 + 2^3 + 3^3 + … + n^3 = n^2(n + 1)^2/4 for all positive integers n

jimthompson5910 (jim_thompson5910):

I'm sure you have the base case already done?

OpenStudy (anonymous):

no. i just started looking at it. i have not tried solving a sequence kind of problem yet.

OpenStudy (anonymous):

n^2(n^2+1)/4

jimthompson5910 (jim_thompson5910):

you'll use induction here

OpenStudy (anonymous):

i have less idea about this problem then i did the last :|

jimthompson5910 (jim_thompson5910):

the base case is n = 1

jimthompson5910 (jim_thompson5910):

prove that is true